Grade 12th passMechanics

A mass of 20kg is falling from a height of 1m and penetrates the sand to 1m. What is the resistance force of the sand?

AKASH KUKREJA
7 Years agoGrade 12th pass

1 Answer

Arun
7 Years ago
Velocity just before impact is given by: 

vf^2 = vi^2 + 2as 
vf^2 = 0 + 2 x 9.81 x 1 = 19.62 
vf = 4.429 m/s 

Deceleration in sand is found from: 
vf^2 = vi^2 + 2as 
0 = 4.429^2 + 2 x a x 1 
-2a = 19.62 
a = - 9.81 

Force = mass x acceleration = 20 x 9.81 = 196.2 N